OpenStudy (anonymous):

Hi! I need to find the values of x, y, and z. I'm not really looking for just the correct answer, but HOW to solve it. x-2y+3z=3 2x+y+5z=8 3x-y-3z=-22

OpenStudy (johnweldon1993):

So we are solving a system of 3 equations...same techniques used for a system of 2 equations \[\large x-2y + 3z = 3\] \[\large 2x + y + 5z = 8\] \[\large 3x - y - 3z = -22\] Now we can either eliminate one of these equations...or we can do a BUNCH of substitutions...your call :D

OpenStudy (johnweldon1993):

Ehh, I'll show the substitution...either way is fine though \[\left(\begin{matrix} x - 2y + 3z = 3 \\ 2x + y + 5z = 8 \\ 3x - y - 3z = -22 \end{matrix}\right)\] So lets solve that first equation for 'x' \[\large x = 3 + 2y - 3z\] Sub that into the second and third equations so now all 3 we have are \[\left(\begin{matrix}x = 3 + 2y - 3z \\ 5y - z = 2 \\5y - 12z = -31 \end{matrix}\right)\] Now solve that second equation for 'y' \[\large y = \frac{2 + z}{5}\] Plug that into the third equation for 'y' The 3 equations we have now are \[\left(\begin{matrix}x = 3 + 2y - 3z \\ y = \frac{2 + z}{5} \\-11z = -33 \end{matrix}\right)\] It can now be seen, that last equation we have...can be solved...we would find z = 3 You can then use that in equation 2...to find y = 1 And finally knowing both we can use equation 1 to find x = -4

OpenStudy (anonymous):

This was super helpful :) Thank you!